OpenStudy (anonymous):

from a barrel of colored marbles you randomly select 1 blue 6 yellow 7 red 4 green and 6 purple marbles find the experimental probability of randomly selecting a marble that is yellow or red

OpenStudy (anonymous):

@jeanoboyle1 do you mean that the barrel contains 1 blue 6 yellow 7 red 4 green and 6 purple ?

OpenStudy (anonymous):

assuming it to be so, the probability is P=6/24+7/24=13/24